OpenStudy (anonymous):

hey! quick question.... how does implicit differentiation help simply work in a related rates problem?

OpenStudy (kira_yamato):

dy/dt = dy/dx . dx/dt

OpenStudy (kira_yamato):

So for example you are given y in terms of x and you are asked to find rate of change of y at a particular time, given rate of change of x that particular time, you can use the above relationship (implicit differentiation) to form an equation dy/dt = f(x,y,t) and just plug in the value for dx/dt, x, y and t to find the rate of change of y.

OpenStudy (anonymous):

suppose boat A is moving horizontally at 6m/s east and boat B was moving towards the south at some speed. find that speed if after 10 seconds boat A is 100 meters from boat B; you might come up with the following equation; x^2+y^2=s x^2+y^2=100 where x is the horizontal distance and y is the verticle distance implicitly differentiating gives 2xx'+2yy' = 0 2x+2y(dy/dx) = 0 {x' = dx/dx} -x/y = dy/dx after 10 seconds x= 6*10 = 60meters and y = sqrt(10000-3600) = 80meters -60/80 = dy/dx or you could use a ladder

OpenStudy (anonymous):

thanks a lot!